OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 4, 12, 36, 108,

OpenStudy (anonymous):

converges i think your multypliyng 3

OpenStudy (anonymous):

Converges; 484 Diverges Converges; 52 Converges; 160

OpenStudy (anonymous):

are the answers they give me

OpenStudy (psymon):

Well, this can be written as: \[\sum_{i=1}^{\infty} 4(3)^{n}\]

OpenStudy (anonymous):

thanks

OpenStudy (psymon):

This is a geo-series of the form ar^n. When the absolute value of r is greater than or equal to 1, the series diverges. So this is a divergent series.

OpenStudy (anonymous):

where did you get the equation

OpenStudy (anonymous):

o got the answer to be (c) is the right

OpenStudy (anonymous):

i got the answer to be (c) is this right

OpenStudy (psymon):

It is a general form of a geometric series. If you can put the series into a form of ar^n, you can write it in the way I did.

OpenStudy (anonymous):

ahh ok well thank you and yes i do believe that c is right

OpenStudy (anonymous):

so is my answer right???

OpenStudy (anonymous):

yes its right

OpenStudy (psymon):

I would say diverges, but oh well.

OpenStudy (anonymous):

no when i just did the equation it put me into the converges answer of c

OpenStudy (anonymous):

okay so (c) is the right answer

OpenStudy (psymon):

Can you show me how then?

OpenStudy (psymon):

I would like to see your reasoning.

OpenStudy (anonymous):

So would I

OpenStudy (anonymous):

yea if i can find the paper i did it on i will show but i want to see @TNGs reasoning

OpenStudy (anonymous):

i dont really have a reasoning i just use the eqaution and i dont really know how to use this software yet to show you all my work in a decent way

OpenStudy (psymon):

There is no work needed with this problem. There is a general form with specific conditions for convergence or divergence. \[\sum_{i=1}^{\infty} ar ^{n}\] If the absolute value of r is greater than or equal to 1, the series diverges. In this case, your a is 4 and your r is 3. This equation represents that for every subsequent part of the series, you are multiply 4 by an increasingly higher power of 3. Regardless of the formula and its conditions, think of it. When n = 1 you have 4, n = 2 you have 36, n = 3 you have 108, n = 4 you have 324, n = 5 you have 972. This number will get infinitely higher.

OpenStudy (anonymous):

ok i stand corrected i messed up in the equation sorry it was a fools mistake

OpenStudy (anonymous):

jsut hop in so is this problem a diverges

OpenStudy (psymon):

This diverges, correct.

OpenStudy (psymon):

Now just for the sake of finishing the explanation of a geometric series I would want to show the conditions for convergence and how you find the limit. With permission of course.

OpenStudy (anonymous):

go right ahead

OpenStudy (psymon):

Alrighty. Now, if the absolute value of r is between 0 and 1, the series converges. This is the formula to find out what numberit converges to. \[\frac{ a }{ 1-r }\] For example, say you had the series of 108, 36, 12, 4, 4/3......this would represent a series in the form of \[\sum_{i=1}^{\infty} 108(\frac{ 1 }{ 3 })^{n}\] This means my a = 108 and my r = 1/3. Therefore this fits the conditions of convergence and the limit would be \[\frac{ 108 }{ 1-\frac{ 1 }{ 3 } }\] \[\frac{ 108 }{ \frac{ 2 }{ 3 } } = 162\]